OpenStudy (anonymous):

Please simplify each radical, and then simplify the expression by combining like radicals. 3√18 + 2√27 + √8 + √81

OpenStudy (akashdeepdeb):

This is how you have to solve these kinds of questions. \[ 3√18 + 2√27 + √8 + √81\] Break them up in their specific roots. \[ 3 \sqrt{9*2} + 2 \sqrt{3*3*3} + \sqrt{2*2*2} + \sqrt{9*9}\] Find their square roots \[ 9 \sqrt{2} + 6 \sqrt{3} + 2 \sqrt{2} + 9\] Combine like terms \[ 11 \sqrt{2} + 6 \sqrt{3} + 9\] Understood? :)

OpenStudy (anonymous):

got it @AkashdeepDeb